OpenStudy (darkigloo):

What is the area of the largest rectangle that can be inscribed in the ellipse 4x^2+9y^2=36?

OpenStudy (turingtest):

have you made any progress yet?

OpenStudy (darkigloo):

All I know is A=4xy and x=sqrt(36-9y^2/4) I don't know what to do now.

OpenStudy (turingtest):

where do you get A=4xy from?

OpenStudy (turingtest):

oh I see

OpenStudy (turingtest):

right, so substitute for x into A

OpenStudy (turingtest):

if you want the hard way, that is...

OpenStudy (anonymous):

Well, if looks like what you can do is let \(A = f(y)\). Then maximize \(f(y)\).

OpenStudy (anonymous):

By finding \(dA/dy = 0\).

OpenStudy (darkigloo):

So I don't use A=4xy?

OpenStudy (anonymous):

You use it, you substitute x with the other equation for x. This makes it entirely a function of y.

OpenStudy (darkigloo):

Ok so A=4(sqrt(36-9y^2/4))(y)

OpenStudy (darkigloo):

I'm not sure what to do now.

zepdrix (zepdrix):

So you've written your area function in terms of one variable so far? Ok good. So you've successfully applied your `constraint` (the ellipse) to your area function. Now we want to maximize the area function with respect to y. Remember how to find max/min?

OpenStudy (darkigloo):

Nope. That's my problem. I don't remember how to do these kinds of problems. :/

zepdrix (zepdrix):

To find max/min of our area function we'll need to first find `critical points` of the function. Then we'll be able to do some quick tests to classify those points. Mmm before we do that though, I think maybe your x is wrong. Lemme check.

zepdrix (zepdrix):

\[\Large\bf\sf 4x^2+9y^2=36\qquad\to\qquad x=\sqrt{\color{red}{9}-\frac{9}{4}y^2}\]I think you forgot to divide the 36 by 4 as well, yes? :o Oh oh maybe you didn't. Maybe it's the way you wrote it. Use brackets to be clear XD A=4(sqrt([36-9y^2]/4))y Mmm so many brackets -_-

zepdrix (zepdrix):

So we want to take the derivative of our function, then set that derivative equal to zero. Solving for those y's will give us critical points of our y variable.

zepdrix (zepdrix):

So ummm.... this one is a little messy. We'll need to apply the product rule.

zepdrix (zepdrix):

\[\Large\bf\sf A=4y\sqrt{9-\frac{9}{4}y^2}\] \[\Large\bf\sf A'\quad=\quad \color{royalblue}{(4y)'}\sqrt{9-\frac{9}{4}y^2}+4y\color{royalblue}{\left(\sqrt{9-\frac{9}{4}y^2}\right)'}\]

zepdrix (zepdrix):

We need to take the derivatives of these blue parts. This process looks familiar I hope? :o

OpenStudy (darkigloo):

Yes

OpenStudy (darkigloo):

would that be \[(4y')(\sqrt{9-(9/4)y^2})+(4y)(((18/4)y')/2\sqrt{9-(9/4)y^2}) \]

zepdrix (zepdrix):

\[\Large\bf\sf A'\quad=\quad \color{royalblue}{(4y)'}\sqrt{9-\frac{9}{4}y^2}+4y\color{royalblue}{\left(\sqrt{9-\frac{9}{4}y^2}\right)'}\] \[\Large\bf\sf A'\quad=\quad \color{orangered}{(4)}\sqrt{9-\frac{9}{4}y^2}+4y\color{orangered}{\left(\frac{-\dfrac{18}{4}y}{2\sqrt{9-\frac{9}{4}y^2}}\right)}\]Umm close, yes.

zepdrix (zepdrix):

We're taking our derivative `with respect to y`, so we shouldn't end up with any `y'` terms.

zepdrix (zepdrix):

That y in the first part will turn into a 1. The y^2 in the second part will turn into a y, right?

OpenStudy (darkigloo):

Yes. I understand.

zepdrix (zepdrix):

Then we set our first derivative equal to zero, and solve for y.

OpenStudy (darkigloo):

Am I on the right track? \[8\sqrt{9-(9/4)y^2}+18y^2=0\]

zepdrix (zepdrix):

I think your second term is -18y^2, yes?

zepdrix (zepdrix):

We picked up a negative from the derivative of the square root part.

OpenStudy (darkigloo):

Oh ok

OpenStudy (darkigloo):

should I do \[8\sqrt{9-(9/4)y^2}/18=y^2\]

zepdrix (zepdrix):

Oh and also umm.. If you multiplied through by that denominator, shouldn't your first term become, \(\Large\bf\sf 8(9-(9/4)y^2)\) ? Without the root?

OpenStudy (darkigloo):

Ohh right..

zepdrix (zepdrix):

After you distribute the 8, it should work out a little nicer.

OpenStudy (darkigloo):

Does y= sqrt(2) ?

zepdrix (zepdrix):

Hmm that looks right. I got something different when I did it. Lemme check again real quick D:

zepdrix (zepdrix):

Ya sqrt2 looks good! We should get two solutions since y is squared. We actually get \(\Large\bf\sf y=\pm sqrt2\) But -sqrt2 doesn't make sense for a rectangles dimensions so we can ignore that value.

OpenStudy (darkigloo):

oh ok. now do I plug in y to x=sqrt(9-(9/4)y^2)?

zepdrix (zepdrix):

Ya that's a good idea. We've found the y coordinate which maximizes area. That will give us the corresponding x value.

OpenStudy (darkigloo):

so x= sqrt(9/2)?

zepdrix (zepdrix):

Ok good. so now how do you find that maximized area given x and y?

OpenStudy (darkigloo):

plug it into 4xy?

zepdrix (zepdrix):

yes

OpenStudy (darkigloo):

so the answer is 12?

zepdrix (zepdrix):

Mmm yes that seems right! Good job! \c:/

OpenStudy (darkigloo):

Thank you so much!!